OpenStudy (anonymous):

if x=100, find the value of Squareroot(x/16-x/25)

OpenStudy (anonymous):

\[\sqrt{100/16 - 100/25}\] \[\sqrt{100/16 - 4}\] \[\sqrt{9/4}\] \[\sqrt{9/4} = (9/4)^{1/2} = 3/2\]